# MedTransformer: Accurate Alzheimer’s Disease Diagnosis for 3D MRI Images through 2D Vision Transformers

### 12 Jan 2024 v1

@@ Abstract
Automated diagnosis of Alzheimer’s Disease (AD) in brain images is becoming a clinically important technique to support precision and efficient diagnosis and treatment planning. A few efforts have been made to automatically diagnose AD in magnetic resonance imaging (MRI) using three-dimensional CNNs. However, due to the complexity of 3D models, the performance is still unsatisfactory, both in terms of accuracy and efficiency. To overcome the complexities of 3D images and 3D models, in this study, we aim to attack this problem with 2D vision Transformers. We propose a 2D transformer-based medical image model with various transformer attention encoders to diagnose Alzheimer’s disease (AD) in 3D MRI images, by cutting the 3D images into multiple 2D slices. The model consists of four main components: shared encoders across three dimensions, dimension-specific encoders, attention across images from the same dimension, and attention across three dimensions. It is used to obtain attention relationships among multiple sequences from different dimensions (axial, coronal, and sagittal) and multiple slices. We also propose morphology augmentation, an erosion and dilation based method to increase the structural difference between AD and normal images. In this experiment, we use multiple datasets from ADNI, AIBL, MIRAID, OASIS to show the performance of our model. Our proposed MedTransformer demonstrates a strong ability in diagnosing AD. These results demonstrate the effectiveness of MedTransformer in learning from 3D data using a much smaller model and its capability to generalize among different medical tasks, which provides a possibility to help doctors diagnose AD in a simpler way. 11footnotetext: Corresponding author
@@ 1 Introduction
Figure 1: The overall framework of our proposed 2D ViT-Based Model for 3D Medical Imaging (MedTransformer).
Alzheimer’s disease (AD) is a highly common neurodegenerative disorder that is usually diagnosed by structural alterations of the brain mass. Therefore, investigating the structural changes in the brain is usually a critical step in diagnosing AD. To study the structural changes, the physicians usually need to read slices of brain images, typically magnetic resonance imaging (MRI), from the patients and study the structural alterations.
However, the analysis of large and intricate MRI images and the manual extraction of crucial information can pose challenges for physicians. Furthermore, the time-consuming nature of manually analyzing brain 3D MRI, coupled with potential inter- or intra-operator variability issues, increases the vulnerability to misdiagnosis [10]. Therefore, the community has been using machine intelligence to help physicians diagnose AD diseases [28]. In recent years, Convolutional Neural Networks (CNN) have been established with a dominant performance in the medical imaging field [38, 13], due to their effectiveness in extracting meaningful spatial hierarchical features from complex image data.
At the same time, transformers and self-attention mechanisms [45] have gradually surpassed CNNs, becoming a trend in the deep learning arena. The Vision Transformer [12] first demonstrated the effective application of transformer networks to image processing, achieving impressive results. Subsequently, transformer networks have gained widespread adoption in computer vision tasks, including image classification [6], object detection [5], and segmentation [49]. Additionally, ViT shows performance on par with traditional CNN-based methods for extensive databases, presenting the transformer model as a viable competitor to other advanced techniques. However, its performance is less convincing on smaller datasets compared to CNN models. A primary reason is that due to a lack of inductive bias of locality, lower layers of ViT can not learn the local relations well, leading to the representation being unreliable [57]. Moreover, in 3D medical imaging, the scarcity of datasets, largely due to ethical considerations that restrict access [39, 41], costly annotations [54, 50], class imbalance challenges [52], and the significant computational demands of processing high-dimensional data [42], is a notable issue. As a result, purely transformer-based methods have yet to see widespread use in 3D medical image diagnosis.
In addition, CNNs often struggle with high-dimensional data due to their limited receptive field [31, 12]. On the other hand, transformer models, through their self-attention mechanisms, analyze relationships between all elements in the input sequence simultaneously. This capability enables transformers to effectively capture global context and dependencies, proving advantageous in tasks requiring long-range interaction analysis. Such spatial relationships are crucial in 3D MRI images for Alzheimer’s diagnosis [26], where understanding cross-sectional interdependencies is the key.
However, vision transformers are facing new challenges when the data are 3D images. With increased spatial dimensions, the self-attention mechanism demands significantly more computational resources, which can lead to suboptimal outcomes. Some works, such as Axial-DeepLab [47], address this by factorizing 2D attention into two 1D attentions along the height and width axes, sequentially improving results. Moreover, the challenge of overfitting is exacerbated in 3D MRI image analysis. This is primarily due to 3D models having considerably more parameters than their 2D counterparts, posing difficulties in training with the limited medical data available.
@@
Inspired by the problems mentioned above, we propose MedTransformer, a pure transformer-based model that factorizes 3D MRI images into three 2D sequences of slices along axial, coronal and sagittal dimensions. Our goal is to classify Alzheimer’s disease (AD) and normal states in 3D MRI images. The overall architecture of the proposed MedTransformer model is shown in Fig 1. Similar to [27], we also combine multi 2D slices as input and use transformer models to classify. However, unlike their model, which uses 3D+2D CNNs to extract the features among the same dimension, we directly use 2D separate transformer encoders. At the same time, we also build attention encoders across slices from the same dimension and the attention encoders across three dimensions. These encoders can help to efficiently combine the feature information better than just keep training using the slices altogether. Our contribution is as follows:
  • We proposed a new transformer-based architecture to fully combine the feature information among multiple slices and multiple planes. These encoders can maximize the receptive field and combine the features thoroughly and capture the long-range relationship in 3D MRI images. This results in state-of-the-art performance among all the baselines.
  • Inspired by [25, 47], we cut the 3D MRI images into different sequences along axial, coronal, and sagittal dimensions. Each sequence consists of multiple slices from that axis. We use encoder blocks to extract and combine the information. In this way, MedTransformer maintains the same ability to extract the whole features and relationships of 3D MRI images as 3D models, but with less model size which results in a more reliable performance.
@@
  • We proposed a novel cross-attention mechanism and a novel guide patch embedding layer, which can better combine the information between slices and sequences. The cross-attention mechanism can help to combine the feature information more efficiently, and the guide patch embedding layer can provide extra global pattern information to each slice.
@@
  • Considering the structure and difference between AD and normal MRI images, we designed the morphology augmentation methods to augment the data. These methods can not only widen the difference between AD and normal images, but also can help to augment the mild cognitive impairment images, a kind of image representing the prodromal stage of AD.
@@ 2 Related Works
The recent success of the transformer architecture in natural language processing [45] has garnered significant attention in the computer vision domain. The transformer has emerged as a substitute for traditional convolution operators, owing to its capacity to capture long-range dependencies. Vision Transformer (ViT) [12] introduces transformer architecture into the computer vision field and starts a craze in combining transformers and images together. Many works have demonstrated remarkable achievements across various tasks, with several cutting-edge methods incorporating transformers for enhanced learning.
Numerous attention-based methods have been proposed for 3D image classification. Point-BERT [55] first partitions the input point cloud into point patches, and uses a miniPointnet [37] to generate a sequence of point embeddings. 3DMedPT [53] embeds local point cloud context by downsampling the points and then grouping local features similar to DGCNN [48]. At the same time, many existing works also deal with 3D object detection problems. Pointformer [34] captures and aggregates local and global features together to do both indoor and outdoor object detection. 3DERT [32] proposes an encoder-decoder module that can be applied directly on the point cloud for extracting feature information, and then predicting 3D bounding boxes. M3DETR [20] aggregates information from raw points, voxels, and bird-eye view, under a unified transformer-based architecture which enables interactions among multi-representation, multi-scale, multi-location feature attention. Also, image segmentation is a hot topic in the medical imaging field, especially in 3D. UNETR [23] employs a strategy of dividing the input 3D MRI images into non-overlapping patches of uniform size. These patches are subsequently projected into an embedding space using a linear layer. By applying a transformer to the projected patches, UNETR [23] learns sequence representations of the input volume (encoder) and effectively captures global multi-scale information. At the same time, Swin UNETR [22] projects multi-modal input data into a 1D sequence of embedding and uses it as input to an encoder composed of a hierarchical Swin Transformer [29].
Key Differences: These models are all using 3D architecture to deal with 3D input, which is insufficient in medical field due to the high value of medical images and limited dataset size. M3T [27] uses a multi-slice and multi-plane transformer. However, they also use 3D CNN for better feature extraction. At the same time, they only use a traditional transformer encoder to extract attention on all slices from axial, coronal, and sagittal dimensions directly. Unlike them, our 2D MedTransformer utilizes different blocks to first extract features among different slices and dimensions, then use a cross-attention mechanism to combine these features together, which can better release the abilities of transformer architecture.
@@ 2.2 Deep Learning for Medical Image Analysis
With the success of deep learning models, extensive research interest has been devoted to deep learning for the development of novel medical image processing algorithms, resulting in remarkably successful deep learning-based models that effectively support disease detection and diagnosis in various medical imaging tasks [9]. U-Net and its variants dominate medical image analysis, which is widely used in image segmentation. Attention U-Net [33] incorporates attention gates into the U-Net architecture to learn important salient features and suppress irrelevant features. Phiseg [3] applys conditional VAE for inference in the U-Net architecture and Uses a separate latent variable to control segmentation at each resolution level to hierarchically generate final segmentations.
@@
For medical image classification, AG-CNN [19] uses the attention mechanism to identify discriminative regions from the global image and fuse the global and local information together to better diagnose thorax disease from chest X-rays. MedicalNet [8] uses the resnet-based [24] model with transfer learning to solve the problem of lacking datasets.  [7] proposes a new self-supervised pretext task based on context restoration to learn high-quality features from unlabeled images.  [40] cuts 3D images into 2D slices and uses paralleled CNNs for each slice with KNN, SVM, and Random forest for classification.
@@
Key Differences: These methods usually focus on CNN based model, which has been outperformed by transformer-based models. In our work, we use a pure transformer-based model to do Alzheimer’s classification and have demonstrated MedTransformer can outperform CNN-based models in classification accuracy results.
3 Methods
Figure 2: The detailed architecture for our MedTransformer.
3.1 An Overview of Vision Transformer
Vision Transformer(ViT) [12] is becoming more and more popular because of its ability to extract long-range features of an image. It reshapes a 2D image x $\in$ $\mathbb{R}^{H\times W\times C}$ into an embedding sequence $\textbf{x}_{p}$in $\mathbb{R}^{N\times(P^{2}\cdot C)}$ with image patch patch $\in$ $\mathbb{R}^{P\times P}$. Because that Transformer uses constant latent vector size D through all of its layers, so ViT flattens the patches and map to D dimensions with a trainable fully connected layer. The output patch embedding is $\textbf{x}_{p}$in $\mathbb{R}^{N\times D}$. Then ViT concatenates a learnable $[class]$ token $\textbf{x}_{class}$in $\mathbb{R}^{1\times D}$ at the first position of the embedding sequence similar to BERT [11] for classification. At the same time, to better learn the position relationship between different image patches, ViT also directly adds a learnable position embedding $\textbf{E}_{pos}$in $\mathbb{R}^{(N+1)\times D}$ into the input sequence for learning the relationship.
$\textbf{S}_{0}$ $=\textbf{x}_{class}\oplus\textbf{x}_{p}+\textbf{E}_{pos}\quad\quad$ $\textbf{S}\in\mathbb{R}^{(N+1)\times D}$ (1)
$\textbf{S}_{l}^{\prime}$ $={\rm MSA(LN(}\textbf{S}_{l-1}))+\textbf{S}_{l-1}$ $l=1...L$ (2)
$\textbf{S}_{l}$ $={\rm MLP(LN(}\textbf{S}_{l}^{\prime}))+\textbf{S}_{l}^{\prime}$ $l=1...L$ (3)
out $={\rm LN(}(\textbf{x}_{class})_{L})$ (4)
label $={\rm MLP}(\textbf{out})$ (5)
The Transformer Encoder has $L$ layers of multiheaded self-attention (MSA) and MLP blocks. At the same time, Layer Norm (LN) and residual connections are applied before and after blocks.
3.2 MedTransformer Architecture
To better utilize the attention ability but avoid the weakness that ViT is highly sensitive to position information, we propose MedTransformer: a pure transformer-based 2D architecture that focuses on multiple slices and multiple views of 3D images. Figure 2 shows the details. MedTransformer mainly consists of 4 blocks:
@@ 3.2 MedTransformer Architecture
First, we cut the 3D image into multiple slices along different views. To better obtain the complete information of the 3D image, we cut each image along three views: sagittal view, coronal view, and axial view. We use $n$ images from each view as the model input. Then similar to ViT, MedTransformer also uses the image patch and patch embedding method to embed the 2D images into 3 sequences including $3\times n$ slices with guide patch embedding layer $\textbf{x}_{guide}$, then concatenates them together as the input to the transformer encoders (Eq. 6). The guide patch embedding aims to reshape the whole sequence into a sequence of flattened 2D patches that has the same shape as the sequence after the normal patch, which means the guide patch embedding has the input channel with the number $3\times n$. By doing this, we can simply add the global information into each special slice sequence. MedTransformer adds learnable position embedding to $\textbf{S}_{0}$ for the patch embedding sequence. (Eq. 7).
@@
Second, the lower layer encoders take $\textbf{S}_{0}$ as input and learn the bias attention among multiple slices and multiple views. To be more specific, the shared Self-Attention Encoders (SAE) across three view dimensions are designed to learn not only the attention of the slice itself but also the relationship between all slices. These networks are Siamese networks [21] which share the same weights. Also, we set the $[class]$ token of each slice $\textbf{x}_{class}^{s}=\textbf{x}_{class}$ before training the network.
@@
The Dimension-specific Self-Attention Encoders (DS-AE) also aim to learn the attention of the slice itself. However, compared with SAE, these encoders only learn the relationship between the slices from the same dimension sequence. So in MedTransformer design, there will be three different encoders, each having the same number of layers $L_{\textnormal{DSAE}}$. In the following equation, t means the three different views.
@@
Then the output of DSAE from each view will be the input of Intra-dimension Cross-Attention Encoders (IntraCAE) of each view. Here MedTransformer will apply cross embedding mechanism to the input embeddings. (Details are in section 3.3.) After the IntraCAE, the embeddings will combine the features from different slices of the same view sufficiently. Following the architecture of DSAE, each view-dependent encoder will have $L_{\textnormal{IntraCAE}}$ layers.
@@
After combining the features between slices of the same dimension independently, the Inter-dimension Cross-Attention Encoders (InterCAE) are proposed to learn the inter-dimension relationship between different sequences from different views. InterCAE will apply cross embedding mechanism again into the view-dependent embeddings. Here MedTransformer only has one encoder with $L_{\textnormal{InterCAE}}$ layers.
@@
Finally, the $[class]$ tokens of the output from three dimensions will be averaged and sent to Layer Norm and classification MLP head as Eq. 4 and Eq. 5 to get the final diagnosis result: AD or normal.

### 29 Feb 2024 v2

# ADAPT: Alzheimer’s Diagnosis through Adaptive Profiling Transformers

@@ Abstract
Automated diagnosis of Alzheimer’s Disease (AD) from brain imaging, such as magnetic resonance imaging (MRI), has become increasingly important and has attracted the community to contribute many deep learning methods. However, many of these methods are facing a trade-off that 3D models tend to be complicated while 2D models cannot capture the full 3D intricacies from the data. In this paper, we introduce a new model structure for diagnosing AD, and it can complete with 3D model’s performances while essentially is a 2D method (thus computationally efficient). While the core idea lies in new perspective of cutting the 3D images into multiple 2D slices from three dimensions, we introduce multiple components that can further benefit the model in this new perspective, including adaptively selecting the number of sclices in each dimension, and the new attention mechanism. In addition, we also introduce a morphology augmentation, which also barely introduces new computational loads, but can help improve the diagnosis performances due to its alignment to the pathology of AD. We name our method ADAPT, which stands for Alzheimer’s Diagnosis through Adaptive Profiling Transformers. We test our model from a practical perspective (the testing domains do not appear in the training one): the diagnosis accuracy favors our ADAPT, while ADAPT uses less parameters than most 3D models use.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: The overall framework of our proposed 2D ViT-Based Model for 3D Medical Imaging (ADAPT).
Alzheimer’s disease (AD) is a highly common neurodegenerative disorder that is usually diagnosed by structural alterations of the brain mass. Assessing an AD usually involves the acquisition of magnetic resonance imaging (MRI) images, since it offers accurate visualization of the anatomy and pathology of the brain (Zhou et al., 2023b). To overcome the vulnerability of misdiagnosis (Despotović et al., 2015) and to speed the diagnosis process, the community has been using machine intelligence to help physicians diagnose AD diseases (Jo et al., 2019).
Considering the complex structure of brain magnetic resonance imaging (MRI), in recent years, Convolutional Neural Networks (CNNs) have been established with a dominant performance in the AD-related field (Salehi et al., 2020; Farooq et al., 2017), due to their effectiveness in extracting meaningful spatial hierarchical features from complex images. Many methods (Zhu et al., 2021; Wen et al., 2020) try to learn the characteristics of AD using CNN-based models. However, the original MRI is complex 3D data, with the proposed 3D model, the input of the 3D convolution operation introduces a third dimension, which greatly increases the burden on the computer. So they use a bag of patches selected from the skull-stripped brain region. These approaches disregard the global context information, which can have a substantial impact on accurately identifying lesions during inference (Wang et al., 2022). Moreover, CNNs are not well-suited for mining global long-dependent information due to their inherent focus on extracting local information (Luo et al., 2016; Dosovitskiy et al., 2020).
Transformers (Vaswani et al., 2017) have also been widely used in medical imaging because of their superior performances over CNNs. Such spatial relationships are crucial in 3D MRI images for Alzheimer’s diagnosis (Iaccarino et al., 2021), where understanding cross-sectional interdependencies is the key. However, transformer-based methods have yet to see widespread use in 3D medical image diagnosis. A primary reason is that due to a lack of inductive bias of locality, lower layers of ViT can not learn the local relations well, leading to the representation being unreliable (Zhu et al., 2023). Moreover, in 3D medical imaging, the scarcity of datasets, largely due to ethical considerations that restrict access (Setio et al., 2017; Simpson et al., 2019), costly annotations (Yu et al., 2019; Wang et al., 2023), class imbalance challenges (Yan et al., 2019), and the significant computational demands of processing high-dimensional data (Tajbakhsh et al., 2020), is a notable issue.
At the same time, these models typically treat all the dimensions in the same way. In contrast, when physicians read the MRI, they usually pay different attentions to different dimensions of the images, according to the atrophic patterns of the brain. This adaptive strategy of the physicians allows them to diagnose more efficiently and accurately.
@@
Inspired by the above, we propose ADAPT, a pure transformer-based model that leverages the captured different features from each view dimension more smartly and efficiently. Our goal is to classify Alzheimer’s disease (AD) and normal states in 3D MRI images. The overall architecture of the proposed ADAPT model is shown in Fig 1. ADAPT factorizes 3D MRI images into three 2D sequences of slices along axial, coronal and sagittal dimensions. Then we combine multiple 2D slices as input and use a 2D separate transformer encoder model to classify. At the same time, we also build attention encoders across slices from the same dimension and the attention encoders across three dimensions. These encoders can help to efficiently combine the feature information better than just keep training using the slices altogether. Benefiting from the special encoder blocks, ADAPT can learn the AD pathology just using a few slices instead of inputting all 2D images, which can further reduce memory footprint. Our contribution is as follows:
  • We proposed a new transformer-based architecture to solve the real-world AD diagnosis problem.
@@
  • We proposed a novel cross-attention mechanism and a novel guide patch embedding, which can gather the information between slices and sequences better.
@@
  • Considering the structure and difference between AD and normal MRI images, we designed the morphology augmentation methods to augment the data.
  • We proposed an adaptive training strategy in order to guide the attention of our model, leading the model to adaptively pay more attention to the more important dimension.
  • Overall, we name our method ADAPT, which is evaluated as the state-of-the-art performance among all the baselines while occupying minimum memory.
@@ 2 Related Works
The recent success of the transformer architecture in natural language processing (Vaswani et al., 2017) has garnered significant attention in the computer vision domain. The transformer has emerged as a substitute for traditional convolution operators, owing to its capacity to capture long-range dependencies. Vision Transformer (ViT) (Dosovitskiy et al., 2020) introduces transformer architecture into the computer vision field and starts a craze in combining transformers and images together. Many works have demonstrated remarkable achievements across various tasks, with several cutting-edge methods incorporating transformers for enhanced learning.
Some attention-based methods have been proposed for 3D image classification. COVID-VIT (Zhou et al., 2023a) uses 3D vision transformers to exploit CT chest information for the accurate classification of COVID. I3D (Carreira & Zisserman, 2017) proposes a new two-stream inflated 3D ConvNet to learn seamless spatio-temporal feature extractors from video, which can be used to do human action classification. At the same time, many existing works also deal with 3D object detection problems. Pointformer (Pan et al., 2021) captures and aggregates local and global features together to do both indoor and outdoor object detection. 3DERT (Misra et al., 2021) proposes an encoder-decoder module that can be applied directly on the point cloud for extracting feature information, and then predicting 3D bounding boxes. Also, image segmentation is a hot topic in the both computer vision and medical imaging fields. Swin UNETR (Hatamizadeh et al., 2021) projects multi-modal input data into a 1D sequence of embedding and uses it as input to an encoder composed of a hierarchical Swin Transformer (Liu et al., 2021).
Key Differences: These models are all using 3D architecture to deal with 3D input, which is inefficient in medical field due to the high value of medical images and limited dataset size. Unlike them, our 2D ADAPT utilizes different blocks to first extract features among different slices and dimensions, then use a cross-attention mechanism to combine these features together, which can better release the abilities of transformer architecture.
@@ 2.2 Deep Learning for Medical Image Analysis
With the success of deep learning models, extensive research interest has been devoted to deep learning for the development of novel medical image processing algorithms, resulting in remarkably successful deep learning-based models that effectively support disease detection and diagnosis in various medical imaging tasks (Chen et al., 2022). U-Net and its variants dominate medical image analysis, which is widely used in image segmentation. Attention U-Net (Oktay et al., 2018) incorporates attention gates into the U-Net architecture to learn important salient features and suppress irrelevant features.
@@
For medical image classification, AG-CNN (Guan et al., 2018) uses the attention mechanism to identify discriminative regions from the global 2D image and fuse the global and local information together to better diagnose thorax disease from chest X-rays. MedicalNet (Chen et al., 2019) uses the resnet-based (He et al., 2016) model with transfer learning to solve the problem of lacking datasets. DomainKnowledge4AD (Zhou et al., 2023b) uses ResNet18 to extract high-dimensional features and proposes domain-knowledge encoding which can capture domain-invariant features and domain-specific features to help predict AD. M3T (Jang & Hwang, 2022) tries to leverage CNNs to capture the local features and use traditional transformer encoders for a long-range relationship in 3D MRI images.
@@
Key Differences: These methods usually focus on CNN based model to extract and combine features, which has been outperformed by transformer-based models. M3T tries to concate transformer blocks after CNNs, however, they propose a much bigger model and treat all slices as the same which is inefficient. In our work, we use a pure transformer-based model with different kinds of encoders to do Alzheimer’s classification and have demonstrated ADAPT can outperform other deep learning models in both classification accuracy results and model size.
3 Methodology
Figure 2: The detailed architecture for our ADAPT. ADAPT consists of four main components: Self-Attention Encoders (SAE) across three views, Dimension-specific Self-Attention Encoders (DS-AE), Intra-dimension Cross-Attention Encoders (IntraCAE), Inter-dimension Cross-Attention Encoders (InterCAE). The figure shows more details in sagittal view for illustration purposes. In practice, the model will adaptively attend to different views.
3.1 ADAPT Architecture
Our model ADAPT builds upon the ViT architecture, extending it to model 3D images. Figure 2 shows the details. ADAPT mainly consists of 4 blocks:
@@ 3.1 ADAPT Architecture
The design inspired from a real-world setting where physicians will pay different attentions to different views, according to the brain pattern. Thus, we design the encoders so that they not only can extract and fusion features from local and global but also can give different weights to different views. To be specific, first, to better obtain the complete information of the 3D image, we cut each image along three views: sagittal view (along x-axis), coronal view (along y-axis), and axial view (along z-axis). We use $n$ images from each view as the model input. Then similar to ViT, ADAPT also uses the image patch and patch embedding method to embed the 2D images into 3 sequences including $3\times n$ slices with guide patch embedding layer $\textbf{x}_{guide}$, then concatenates them together as the input to the transformer encoders (Eq. 1). The guide patch embedding aims to reshape the whole sequence into a sequence of flattened 2D patches that has the same shape as the sequence after the normal patch, which means the guide patch embedding has the input channel with the number $3\times n$. Thus, we can use 3D models to extract the global information and add it to each special slice sequence. Because our model mainly focuses on 2D slice dimension, guide patch embedding can help to keep the relative position information of 3D brain.
@@
Second, the lower layer encoders learn the bias attention among multiple slices and multiple views. To be more specific, the shared Self-Attention Encoders (SAE) across three view dimensions are designed to learn not only the attention of the slice itself but also the relationship between all slices. The designed encoder can realize global information extraction for the first time. These encoders can also help to keep the relative position information of 3D MRI. These networks are Siamese networks (Guo et al., 2017) which share the same weights.
@@
The Dimension-specific Self-Attention Encoders (DS-AE) also aim to learn the attention of the slice itself. However, compared with SAE, these encoders focus more on the relationship between the slices from the same dimension sequence. These encoders can better extract the local features from the same view dimension. This will fill the gap that transformers cannot capture the local features well however the local embeddings of different brain tissues (such as hippocampus and cortex) are really important in AD diagnosis. In the following equation, t means the three different views.
@@
We will fusion the local features from the same dimension first. So we design Intra-dimension Cross-Attention Encoders (IntraCAE). Here ADAPT will apply cross embedding mechanism to the input embeddings. (Details are in section 3.2.) After the IntraCAE, the embeddings will gather the features from different slices of the same view sufficiently.
@@
After combining the features between slices of the same dimension independently, the last Inter-dimension Cross-Attention Encoders (InterCAE) are proposed to learn the inter-dimension relationship among different sequences from different views. This is corresponding to the SAE layer and will gather the global features together. InterCAE will apply cross embedding mechanism again into the view-dependent embeddings.
@@
Finally, the $[class]$ tokens of the output from three dimensions will be averaged and sent to Layer Norm and classification MLP head as Eq. LABEL:4 and Eq. LABEL:5 to get the final diagnosis result: AD or normal.